hello, can you give me the solution set of ordered pairs to
y=5x-3 and y=x+5 and show how you arrived at the answer?
thank you much
p.s. their answer is (2,7)

Answer by ikleyn(53763)

.
hello, can you give me the solution set of ordered pairs to
y=5x-3 and y=x+5 and show how you arrived at the answer?
thank you much
p.s. their answer is (2,7)
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Since the two equations 

y = 5x-3   and
y = x+5

have identical left sides, their right sides are equal too:

5x-3 = x+5.

I am sure you can easily solve it on your own and find x.

Then find y.

In this way, you will get the answer.